OpenStudy (djbreezy):

Marcus is putting a rectangular fence around his garden to keep the animals from stealing his vegetables. The perimeter of the rectangle is 80 feet. The ratio of the width to the length is 3 to 5. What are the dimensions of the rectangle?

OpenStudy (misti786):

let the dimension is x so 5x, 3x

OpenStudy (misti786):

the perimeter of a rectangle is 2(length+width), where length is 5x and width is 3x), and the equation would be 2(5x+3x)=80, now you can solve the equation and find the x, then substitute the x and you will find the dimension.

OpenStudy (welshfella):

the ratio 3:5 means that the width = 3 / 5 of the length. so if the length is x then the width is (3/5)x = 0.6x

OpenStudy (welshfella):

misti786 ' s method will also work

OpenStudy (djbreezy):

@welshfella ......can u show me how to do it give all the work and the answer so therefore next time i know how to do it next time i end up doing a question like this?

OpenStudy (djbreezy):

or @misti786 ??

OpenStudy (misti786):

so 2(8x)=80 16x=80 divide both side by 16 x=5 Therefore the length is 3*5=15 and width is 5*5=25

OpenStudy (djbreezy):

Thx

OpenStudy (welshfella):

well the length is 25 and width 15 - that was just a typo.

OpenStudy (welshfella):

by definition, length is greater than width.